# KAN4TSF: Are KAN and KAN-based models Effective for Time Series Forecasting?

### v1

@@ 4.1 Mixture of KAN Experts Layer
ETTh1,h2 7 17,420 1 hour
ETTm1,m2 7 69,680 15 min
@@
Table 2: Overall performance. Both the best results of KAN-based models and baselines are highlighted in bold, and the best results of all methods are marked in red.
Dataset P KAN-based Transformer-based CNN-based Linear-based
RMoK-S RMoK-B PatchTST FEDFormer TimesNet SCINet RLinear DLinear
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.326 0.360 0.320 0.358 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.355 0.376 0.345 0.372
192 0.367 0.382 0.364 0.383 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.391 0.392 0.380 0.389
336 0.400 0.404 0.395 0.405 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.424 0.415 0.413 0.413
720 0.462 0.439 0.457 0.440 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.450 0.474 0.453
avg 0.389 0.396 0.384 0.397 0.387 0.400 0.448 0.452 0.400 0.406 0.485 0.481 0.414 0.407 0.403 0.407
ETTm2 96 0.176 0.261 0.176 0.261 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.182 0.265 0.193 0.292
192 0.244 0.305 0.240 0.302 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.246 0.304 0.284 0.362
336 0.306 0.346 0.299 0.342 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.307 0.342 0.369 0.427
720 0.405 0.404 0.397 0.401 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.407 0.398 0.554 0.522
avg 0.283 0.329 0.278 0.326 0.281 0.326 0.305 0.349 0.291 0.333 0.571 0.537 0.286 0.327 0.350 0.401
ETTh1 96 0.382 0.396 0.374 0.397 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.386 0.395 0.386 0.400
192 0.430 0.426 0.419 0.429 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.437 0.424 0.437 0.432
336 0.468 0.443 0.461 0.450 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.479 0.446 0.481 0.459
720 0.450 0.458 0.474 0.467 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.481 0.470 0.519 0.516
avg 0.433 0.431 0.432 0.436 0.469 0.454 0.440 0.460 0.458 0.450 0.747 0.647 0.446 0.434 0.456 0.452
ETTh2 96 0.313 0.357 0.301 0.353 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.288 0.338 0.333 0.387
192 0.397 0.409 0.379 0.405 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.374 0.390 0.477 0.476
336 0.441 0.446 0.432 0.446 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.415 0.426 0.594 0.541
720 0.453 0.461 0.446 0.463 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.420 0.440 0.831 0.657
avg 0.401 0.418 0.390 0.417 0.387 0.407 0.437 0.449 0.414 0.427 0.954 0.723 0.374 0.398 0.559 0.515
ECL 96 0.187 0.273 0.178 0.267 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.201 0.281 0.197 0.282
192 0.193 0.280 0.187 0.274 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.201 0.283 0.196 0.285
336 0.210 0.295 0.204 0.290 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.215 0.298 0.209 0.301
720 0.256 0.330 0.247 0.323 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.257 0.331 0.245 0.333
avg 0.211 0.294 0.204 0.288 0.205 0.290 0.214 0.327 0.192 0.295 0.268 0.365 0.219 0.298 0.212 0.300
Traffic 96 0.601 0.367 0.541 0.340 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.649 0.389 0.650 0.396
192 0.576 0.354 0.529 0.330 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.601 0.366 0.598 0.370
336 0.588 0.359 0.545 0.334 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.609 0.369 0.605 0.373
720 0.625 0.376 0.580 0.351 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.647 0.387 0.645 0.394
avg 0.597 0.364 0.549 0.339 0.481 0.304 0.610 0.376 0.620 0.336 0.804 0.509 0.626 0.378 0.625 0.383
Weather 96 0.175 0.225 0.171 0.221 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.192 0.232 0.196 0.255
192 0.224 0.267 0.220 0.263 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.240 0.271 0.237 0.296
336 0.284 0.308 0.277 0.302 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.292 0.307 0.283 0.335
720 0.366 0.358 0.360 0.354 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.364 0.353 0.345 0.381
avg 0.262 0.290 0.257 0.285 0.259 0.281 0.309 0.360 0.259 0.287 0.292 0.363 0.272 0.291 0.265 0.317
Top-1 Count 31 23 7 12
@@
RMoK-S RMoL-S RWavKAN RTaylorKAN RLinear
@@
Table 4: The performance of KAN within iTransformer. The best results of all methods are highlighted in bold. KAN, JacobKAN, and TaylorKAN are KANs with B-splines functions, Jacobi polynomials, and Taylor polynomials.
Dataset P iTransformer +KAN +JacobKAN +TaylorKAN +MoK (Ours)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.341 0.373 0.689 0.547 0.340 0.373 0.331 0.365 0.329 0.364
192 0.388 0.397 0.707 0.556 0.382 0.395 0.379 0.392 0.376 0.390
336 0.426 0.420 0.720 0.565 0.418 0.419 0.417 0.416 0.412 0.415
720 0.498 0.458 0.746 0.582 0.487 0.456 0.484 0.450 0.478 0.453
avg 0.413 0.412 0.716 0.562 0.407 0.411 0.403 0.406 0.399 0.406
ETTh1 96 0.390 0.404 0.696 0.557 0.401 0.414 0.397 0.409 0.393 0.409
192 0.445 0.436 0.716 0.572 0.454 0.445 0.443 0.436 0.442 0.441
336 0.489 0.457 0.721 0.583 0.499 0.469 0.491 0.461 0.476 0.458
720 0.505 0.487 0.731 0.607 0.521 0.499 0.510 0.488 0.494 0.485
avg 0.457 0.446 0.716 0.580 0.469 0.457 0.460 0.448 0.451 0.448
4.2 Reversible Mixture of KAN Experts Model
We can get a sophisticated KAN-based model by stacking multiple KANs or replacing the linear layers of the existing models with KANs. However, we try to design a simple KAN-based model that is easy to analyze while achieving comparable performance to the most state-of-the-art time series forecasting methods.
Inspired by several successful single-layer methods Li et al. (2023); Zeng et al. (2023), we propose a simple, effective and interpretable KAN-based model, Reversible Mixture of KAN Experts Network (RMoK), which uses RevIN Kim et al. (2022) and single MoK layer, as shown in Figure 2. First, $\text{RevIN}^{+}$ (the normalization operation of RevIN) uses a learnable affine transformation to normalize the input time series of each variable. Then, the MoK layer obtains the prediction results based on the normalized time series features. Finally, the prediction results are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ (the denormalization operation of RevIN).
During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts. Following the previous work Shazeer et al. (2017), we add an load balancing loss function to encourage experts have equal importance. First, we count the weight of experts as loads, and calcuate the square of the coefficient of variation of the load values as additional loss:
$\mathrm{L}_{\mathrm{load-balancing}}=\mathrm{CV}(\mathrm{loads})^{2}.$ (8)
And the total loss function is the sum of the prediction loss and load balancing loss with weight $w_{l}$:
$\mathrm{L}=\mathrm{MSE}(\mathcal{Y},\hat{\mathcal{Y}})+w_{l}\cdot\mathrm{L}_{\mathrm{load-balancing}}.$ (9)
5 Experiments
In this section, we experimentally validate the effectiveness of our proposed KAN-based models, RMoK, on various time series forecasting benchmarks. Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, the impact of integrating the KAN and MoK layers into other network structures (Transformer), and the interpretability of RMoK.
@@ 5.1 Experimental Settings
Dataset.
We conduct extensive experiments on seven widely-used real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al. (2021), ECL, Traffic and Weather Lai et al. (2018), whose statistical information is shown in Table 1. We follow the same data processing operations used in TimesNet Wu et al. (2023), where the training, validation, and testing sets are divided according to chronological order.
Evaluation Metrics
Following previous works Zhou et al. (2021); Wu et al. (2023), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
@@
We select 6 well-acknowledged forecasting models as our baselines, including (1) Transformer-based methods: PatchTST Nie et al. (2023) and FEDformer Zhou et al. (2022); (2) CNN-based methods: TimesNet Wu et al. (2023) and SCINet Liu et al. (2022b); (3) Linear-based methods: RLinear Li et al. (2023) and DLinear Zeng et al. (2023).
Table 5: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1. Training batch size is 64 and infer batch size is 1.
Model Param Train (it/s) Infer (it/s)
@@
RKAN 1.2 M 5.94 193.25
RWKAN 277 K 70.39 308.69
RTKAN 208 K 77.87 316.94
RJKAN (d=4) 345 K 76.77 280.06
RJKAN (d=6) 483 K 58.96 254.65
RMoK-S 972 K 34.86 190.57
5.2 Can KAN-based Models get SOTA Performance?
In this section, we conduct extensive experiments to compare the forecasting performance of KAN-based models with advanced baselines.
We use two versions of our proposed RMoK: RMoK-S (the small version, which has four experts), and RMoK-B (the base version, which has eight experts). We compare RMoKs with six popular Transformer-based, CNN-based and Linear-based baselines on seven benchmarks. The experimental results are shown in Table 2, where both the best results of KAN-based models and baselines are highlighted in bold and the best results of all models are marked in red. The results of baselines come from published papers Wu et al. (2023), and the results of RMoKs are average of four times experiments with fixed seed in [0, 1, 2, 3]. Surprisingly, RMoK achieves the best results in most cases. Considering that RMoK is a simple single-layer method that does not model correlations among variates, this empirical finding adequately demonstrates that KAN-based models are effective in the TSF task.
Specifically, we can divide these seven datasets into two groups based on the number of variables. On the four ETT subsets with fewer variables, RMoK outperforms the baselines most, and RMoK-S and RMoK-B exhibit their own strengths under different forecasting length $P$. On the Weather, ECL, and Traffic datasets with more variables, RMoK-B is significantly better than RMoK-S, indicating that the mixture of experts approach is suitable for dealing with a large number of variables. In addition, due to the intricate spatiotemporal correlation among various variables in Traffic dataset, the Transformer-based method PatchTST achieves the best results, whereas our proposed simple RMoK achieves suboptimal results that far exceeded other baselines.
5.3 Does KAN Outperform Linear?
In this section, we conduct ablation experiments to compare KAN-based models with Linear-based models in three time series forecasting datasets.
For fair comparison, we replace the four KAN experts in RMoK-S with four Linear experts to obtain a Linear-based baseline with the mixture of expert structure, naming RMoL-S. And we replace the entire MoK layer with a single KAN or Linear to obtain RWavKAN, RTaylorKAN and RLinear to analyze the performance of KAN variants and Linear on the TSF task. The experimental results are reported in Table 3, where all the results are average of 4 times experiments and the best results are highlighted in bold. We can conclude three useful empirical experiences. First, the KAN-based model outperforms the linear-based model in most cases. We speculate this phenomenon is due to KAN’s function representation idea, which is more efficient to capture the periodicity and trend in time series. Second, the mixture of expert structures is applicable to both KAN and Linear, which should be attributed to the fact that the gating network assigns variables to specific experts. Third, the performance of KAN-based models is affected by the specific function, which may be related to the intrinsic distribution of the time series.

### v2

@@ 3.4 Training Strategy
ETTh1/h2 7 17,420 1 hour
ETTm1/m2 7 69,680 15 min
@@
Table 1: The statistics of the used datasets.
MMK iTransformer PatchTST FEDformer TimesNet SCINet TiDE DLinear
Models MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.374 0.397 0.386 0.405 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.479 0.464 0.386 0.400
192 0.419 0.429 0.441 0.436 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.525 0.492 0.437 0.432
336 0.461 0.450 0.487 0.458 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.565 0.515 0.481 0.459
ETTh1 720 0.474 0.467 0.503 0.491 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.594 0.558 0.519 0.516
96 0.301 0.353 0.297 0.349 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.400 0.440 0.333 0.387
192 0.379 0.405 0.380 0.400 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.528 0.509 0.477 0.476
336 0.432 0.446 0.428 0.432 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.643 0.571 0.594 0.541
ETTh2 720 0.446 0.463 0.427 0.445 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.874 0.679 0.831 0.657
96 0.320 0.358 0.334 0.368 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.364 0.387 0.345 0.372
192 0.364 0.383 0.377 0.391 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.398 0.404 0.380 0.389
336 0.395 0.405 0.426 0.420 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.428 0.425 0.413 0.413
ETTm1 720 0.457 0.440 0.491 0.459 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.461 0.474 0.453
96 0.176 0.261 0.180 0.264 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.207 0.305 0.193 0.292
192 0.240 0.302 0.250 0.309 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.290 0.364 0.284 0.362
336 0.299 0.342 0.311 0.348 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.377 0.422 0.369 0.427
ETTm2 720 0.397 0.401 0.412 0.407 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.558 0.524 0.554 0.522
96 0.166 0.256 0.148 0.240 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.237 0.329 0.197 0.282
192 0.187 0.274 0.162 0.253 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.236 0.330 0.196 0.285
336 0.204 0.290 0.178 0.269 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.249 0.344 0.209 0.301
ECL 720 0.247 0.323 0.225 0.317 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.284 0.373 0.245 0.333
96 0.511 0.324 0.395 0.268 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.805 0.493 0.650 0.396
192 0.529 0.330 0.417 0.276 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.756 0.474 0.598 0.370
336 0.545 0.334 0.433 0.283 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.762 0.477 0.605 0.373
Traffic 720 0.580 0.351 0.467 0.302 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.719 0.449 0.645 0.394
96 0.164 0.210 0.174 0.214 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.202 0.261 0.196 0.255
192 0.210 0.251 0.221 0.254 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.242 0.298 0.237 0.296
336 0.265 0.290 0.278 0.296 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.287 0.335 0.283 0.335
Weather 720 0.343 0.342 0.358 0.349 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.351 0.386 0.345 0.381
Rank 1 27 21 8 1 2 0 0 0
Table 2: Multivariate time series forecasting results. Forecasting horizons $P$ is 96, 192, 336, 720, and input length $T$ is set as 96. A lower value indicates better performance. The best results are marked in red.
4 Experiments
In this section, we experimentally validate the effectiveness of KAN on seven public benchmarks. Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, integrating the KANs into other advanced models, and the interpretability of MMK.
@@ 4 Experiments
MoK MoL WavKAN TaylorKAN Linear
@@ 4.1 Experimental Settings
Datasets
We conduct extensive experiments on seven widely-used datasets, including ETT(h1, h2, m1, m2) Zhou et al. (2021), ECL, Traffic and Weather Lai et al. (2018), whose statistical information is shown in Table 1. We follow the same data processing operations used in TimesNet Wu et al. (2023), where the training, validation, and testing sets are divided according to chronological order.
Settings
We implement our method and baselines in a unify code library with PyTorch  [^1] https://github.com/pytorch/pytorch and Pytorch-lightning  [^2] https://github.com/Lightning-AI/pytorch-lightning , and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs. In order to prevent the ‘Drop Last’ Qiu et al. (2024) in the test stage where the samples of the last incomplete batch are discarded, we set the test batch size to 1. We use the Mean Squared Error (MSE) and the Mean Absolute Error (MAE) as metrics. We use KAN, TaylorKAN, JacobiKAN and WavKAN to construct the MoK layer. The code in the supplementary files provides more details.
@@
We select seven well-known baselines, including (1) Transformer-based methods: iTransformer Liu et al. (2024a), PatchTST Nie et al. (2023) and FEDformer Zhou et al. (2022); (2) CNN-based methods: TimesNet Wu et al. (2023) and SCINet Liu et al. (2022a); (3) Linear-based methods: TiDE Das et al. and DLinear Zeng et al. (2023).
4.2 Can KANs Get the SOTA Performance?
In this section, we compare our proposed MMK with 7 popular Transformer-based, CNN-based and Linear-based baselines on seven custom datasets to comprehensively evaluate the performance of KAN. The experimental results are shown in Table 2, where the best results of all models are marked in red and the second best results are marked with underlines. The results of baselines come from previous paper Liu et al. (2024a), and the results of MMK are average of four times experiments with fixed seed [0, 1, 2, 3]. Surprisingly, MMK achieves the best performance in most cases. And it significantly outperforms on the Weather dataset, reducing the MSE by 5.74% on the 96-step forecasting compared to the second-best method, iTransformer. We infer that this improvement comes from the fact that KAN’s function-fitting ability is more suitable for processing natural science data. The performance on traffic data is poor because our method does not design a module to exploit the correlation between variables. Subsequent research can consider designing a KAN-based variable relationship mining module to better improve model performance. In general, a properly designed KAN-based model can achieve the state-of-the-art performance in most cases while maintaining the advantages of interpretability.
Dataset iTransformer +KAN +MoK
MSE MAE MSE MAE MSE MAE
ETTh1 96 0.390 0.404 0.397 0.409 0.393 0.409
192 0.445 0.436 0.443 0.436 0.442 0.441
336 0.489 0.457 0.491 0.461 0.476 0.458
720 0.505 0.487 0.510 0.488 0.494 0.485
avg 0.457 0.446 0.460 0.448 0.451 0.448
ETTm1 96 0.341 0.373 0.331 0.365 0.329 0.364
192 0.388 0.397 0.379 0.392 0.376 0.390
336 0.426 0.420 0.417 0.416 0.412 0.415
720 0.498 0.458 0.484 0.450 0.478 0.453
avg 0.413 0.412 0.403 0.406 0.399 0.406
Table 4: The performance of KAN within iTransformer. The best results of all methods are highlighted in red.
4.3 Does KANs Outperform Linear?
In this section, we conduct ablation experiments to compare KAN with Linear in time series forecasting tasks. For fair comprison, we only use one layer of forward network that means all models in this experiment only consist of RevIN and a single network layer.
  • MoK: It uses four KAN variants (KAN, WavKAN, TaylorKAN, and JacobiKAN) as experts for forecasting.
  • MoL: It uses four Linears as experts for forecasting.
  • WavKAN, TaylorKAN, Linear: They use the corresponding layer for forecasting.
The experimental results are reported in Table 3, where all the results are the average of 4 times experiments, and the best results are highlighted in red. We can conclude three useful empirical experiences. First, the KAN-based model outperforms the linear-based model in most cases. We speculate this phenomenon is due to KAN’s function representation idea, which is more efficient to capture the periodicity and trend in time series. Second, the mixture of expert structures is applicable to both KAN and Linear, which should be attributed to the fact that the gating network assigns variables to specific experts. Third, the performance of KAN-based models is affected by the specific function, which may be related to the intrinsic distribution of the time series. Through these two groups of experiments, we can conclude that KAN has competitive performance compared to Linear in time-series forecasting.
Models ETTh1 ETTm1
MSE MAE MSE MAE
TimeMixer 0.375 0.400 0.320 0.357
+MoK 0.375 0.398 0.314 0.353
Improve 0.00% 0.50% 1.88% 1.12%
GBT 0.398 0.418 0.319 0.370
+MoK 0.396 0.417 0.311 0.370
Improve 0.50% 0.24% 2.51% 0.00%
Pathformer 0.382 0.400 0.316 0.346
+MoK 0.379 0.389 0.315 0.343
Improve 0.79% 2.75% 0.32% 0.87%
Table 5: The performance of KAN within other model.
4.4 Can KANs be Integrated into Other Models?
In Section 4.2, we have demonstrated that our proposed MMK has excellent performance compared to models based on other network architectures. In order to further explore the upper bound of KAN’s performance in time series forecasting, we integrate KAN into some existing time series forecasting models in the form of replacing the original predictor so as to make full use of the progress of advanced time series forecasting technology. First, we select iTransformer Liu et al. (2024a) as baseline and replace the last linear predictor with a KAN or MoK layer. For fair comparison, we set the same network hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy. As shown in Table 4, iTransformer with KAN and MoK can achieve better performance in most cases. Then, we intergate MoK with more models Wang et al. (2024); Chen et al. (2024); Shen et al. (2023) in Table 5, MoK can achieve better performance in most cases. These experimental findings demonstrate that MoK is a powerful form for integrating KAN into existing models.
4.5 How to Get Better MMK?
We evaluate the impact of different modules or training strategies on the forecasting performance. The pre-sampling weights initialization strategy is used in all experiments to maintain the stability of training. As shown in Figure 3, we report the MSE of $L$=96 and $P$=96, except for ‘+Longer Input’ where $L$=192 and $P$=96. First, we can find that MoK and RevIN are useful to enhance the forecasting ability. Then, adding model depth (‘+MoK’) only will drop the performance significance. The MoKBlock module and the warmup strategy can improve the performance of multi-layer MMK to achieve the best results. We also report the result under ‘+Longer Input’ to illustrate that MMK can get enhancement from longer input.
Figure 3: The impact of different modules or training strategies.
4.6 Are KANs Efficiency?
We report the size of model parameters, the training and inferring speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720. We set training batch size to 64, infer batch size to 1. As shown in Table 6, MMK has more parameters than DLinear, but significantly less than transformer-based methods like iTransformer and PatchTST. The running speed of MMK is slower than that of DLinear. We believe that as researchers continue to optimize the computational efficiency of KAN, the running speed of MMK will be improved in the future.
Model Param Training (it/s) Infer (it/s)
@@ 4.6 Are KANs Efficiency?
MMK 972 K 34.86 190.57
Table 6: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1.
4.7 Case Study